# Recognising perfect fits

### v1

@@ 5.3   Recognising suspensions
Recall from 4.31 that we have the (multi)slope $\mu({\varphi})$ of the flow meridian(s); this is an element of $H_{1}(\partial M)\simeq\operatorname{\mathbb{Z}}^{2n}$ where $n$ is the number of boundary components of $M^{\circ}$. Observe that the flow $\varphi$ is a suspension flow if only if filling back the original meridian disk caps off a disk in some choice of fibering of $M^{\circ}$. That is, precisely when $\mu({\varphi})$ is a fiber slope for some choice of fiber. To determine when this occurs, we must compute the set of fiber slopes. This means computing the fibered face corresponding to $\mathcal{V}(\varphi)$; doing so has been implemented in the Veering software [Veering]. They compute the rays that bound the cone in $H_{2}(M,\partial M)$. The fiber slopes are the image in $H_{1}(\partial M)\simeq\operatorname{\mathbb{Z}}^{2n}$ of the linear map $H_{2}(M,\partial M)\to H_{1}(\partial M)$. So, this image is also a cone. Then, checking whether $\varphi$ is a suspension amounts to checking if $\mu({\varphi})$ satisfies the corresponding linear inequalities.
@@ 6   Questions
Consider the FindFit routine (Algorithm 3.1). Suppose that $\varphi$ has perfect fits. So, as in the characterisation 2.9, $\varphi$ has a pair of (possibly identical) almost freely homotopic orbits $(\alpha,\beta)$. We define the length of $(\alpha,\beta)$ as the sum of the length of their itineraries. Let $|\mathcal{B}|$ be the number of edges in $\mathcal{M}(\mathcal{B})$.
@@
The closest to a result of this flavour is given by Tsang in [tsang, Theorem 6.4]. They give bounds on the length of orbits needed to be drilled to produce a Birkhoff section, but this requires no perfect fits from the outset.
@@
These questions are closely related to the quasigeodesic properties of the underlying flow. Recall that $\varphi$ is quasigeodesic if there is $q>0$ such that the flow lines in the universal cover are $q$-quasigeodesics. For a general reference, see [calegari-folns-book, §10]. Fenley [fenley-qg, Main Theorem] shows that $\varphi$ is quasigeodesic when it is bounded, which includes when $\varphi$ has no perfect fits. Endow $M$ with a metric by giving boxes height one and aspect ratios following [mosher1996laminations, Lemma 3.1.1]. Following 4.14, knowing the quasigeodesic constant for $\varphi$ would allow us to project $\mathcal{B}$ to $\mathcal{P}$, and allow for a different approach to the FindVeering routine.
@@
[…]

### v2

@@ 5.3  Recognising suspensions
Recall from 4.30 that we have the (multi)slope $\mu({\varphi})$ of the flow meridian(s); this is an element of $H_{1}(\partial M)\simeq\operatorname{\mathbb{Z}}^{2n}$ where $n$ is the number of boundary components of $M^{\circ}$. Observe that the flow $\varphi$ is a suspension flow if only if filling back the original meridian disk caps off a disk in some choice of fibering of $M^{\circ}$. That is, precisely when $\mu({\varphi})$ is a fiber slope for some choice of fiber. To determine when this occurs, we must compute the set of fiber slopes. This means computing the fibered face corresponding to $\mathcal{V}(\varphi)$; doing so has been implemented in the Veering software [Veering]. They compute the rays that bound the cone in $H_{2}(M,\partial M)$. The fiber slopes are the image in $H_{1}(\partial M)\simeq\operatorname{\mathbb{Z}}^{2n}$ of the linear map $H_{2}(M,\partial M)\to H_{1}(\partial M)$. So, this image is also a cone. Then, checking whether $\varphi$ is a suspension amounts to checking if $\mu({\varphi})$ satisfies the corresponding linear inequalities.
@@ 6 Questions
Consider the FindFit routine (Algorithm 3.6). Suppose that $\varphi$ has perfect fits. So, as in the characterisation 2.10, $\varphi$ has a pair of (possibly identical) almost freely homotopic orbits $(\alpha,\beta)$. We define the length of $(\alpha,\beta)$ as the sum of the length of their itineraries. Let $|\mathcal{B}|$ be the number of edges in $\mathcal{M}(\mathcal{B})$.
@@
The closest to a result of this flavour is given by Tsang in [tsang-birkhoff, Theorem 6.4]. They give bounds on the length of orbits needed to be drilled to produce a Birkhoff section, but this requires no perfect fits from the outset.
@@
These questions are closely related to the quasigeodesic properties of the underlying flow. Recall that $\varphi$ is quasigeodesic if there is $q>0$ such that the flow lines in the universal cover are $q$-quasigeodesics. For a general reference, see [calegari-folns-book, §10]. Fenley [fenley-qg, Main Theorem] shows that $\varphi$ is quasigeodesic when it is bounded, which includes when $\varphi$ has no perfect fits. Endow $M$ with a metric by giving boxes height one and aspect ratios following [mosher1996laminations, Lemma 3.1.1]. The metric does not necessarily match across gluings, so we only obtain a path metric; see [agol-tsang, page 3440] for a thorough description. Following 4.13, knowing the quasigeodesic constant for $\varphi$ would allow us to project $\mathcal{B}$ to $\mathcal{P}$, and allow for a different approach to the FindVeering routine.
@@
One can ask the same question when considering orbit equivalence for unmarked flows. We have access to powerful tools through Anosov-like actions [bfm-classi] and veering triangulations (say, in the form of [tsang-friedconjecture, Theorem 6.1]). However, one needs a robust way to compare the nonsingular periodic orbits between flows. Another direction could be the new classifying invariant of geometric types introduced in [geometric-types], where some similar algorithmic problems are considered.